Solve the system of equations by the addition method. If the system does not have a single ordered pair as a? solution, state whether the system is inconsistent or dependent.

5X + 4 = 6

2X + 2Y = 4
Answer #1

Multiply second equation by 2 and subtract this from 1

5x+4y-2*(2x+2y) = 6-4*2

5x+4y-4x-4y=6-8

x = -2

5*(-2)+4y = 6

4y = 16

y = 4

Solution is x = -2 and y = 4